OpenStudy (anonymous):

For a triangle ABC with |AB| = c, |AC| = b, |BC| = a, let M be the midpoint of AB. Find |CM|

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

Say \(\angle AMC = \theta\), then \(\angle BMC = 180-\theta\) as they both are linear pair angles.

ganeshie8 (ganeshie8):

Apply law of cosines in the left side triangle and get \[\large b^2 = |CM|^2+(c/2)^2-2|CM|(c/2)\cos(\theta)\tag{1}\] Apply law of cosines in the right side most triangle and get \[\large a^2 = |CM|^2+(c/2)^2-2|CM|(c/2)\cos(180-\theta)\tag{2}\]

ganeshie8 (ganeshie8):

Adding them gives the desired result but im not so sure if you're allowed to use law of cosines

OpenStudy (anonymous):

I don't think it's not allowed. We were supposed to do it 2 ways. I'm sure something like this is fine for the first way, but for the 2nd way it had to be done with vectors.

ganeshie8 (ganeshie8):